OpenStudy (anonymous):

Enter the equation of the line meeting the given conditions. Please put the equation instandard form. Containing C(3, 3) and D(-3, 2)

OpenStudy (anonymous):

y= 1/6x + 5/2

OpenStudy (anonymous):

or can also rewrite as 6y-x-15=0

OpenStudy (anonymous):

Hi Radar, Great explation except that after cross multiplication you should end up with 6y=x+15 not (6y = x+18) you forgot to subtract the 3.

OpenStudy (radar):

I was trying to figure why we weren't on track.

OpenStudy (radar):

I see it now. I wish I could just go in there to correct it, I want the steps to be there. lol I am gonna try something

OpenStudy (anonymous):

yea, :)

OpenStudy (radar):

calculate slope: m=(3-2)/(3+3)=1/6 1/6=(y-3)/(x-3) cross multiply getting x-3=6y-18 6y=x+18-3 6y=x+15 y=1/6x+15/6 y=1/6x +5/2 this form is in the slope intercept form standard form: y-1/6x-5/2=0

OpenStudy (radar):

I think I had better call it quits for tonight lol.